OpenStudy (anonymous):

ntegrate |x| dx from x=-1 to 1 ans=1 HOw?

OpenStudy (turingtest):

Break up the intergralby remembering the definition of absolue value:\[\left| x \right|=x\] if x>0 \[\left| x \right|=-x\] if x<0 \[\int\limits_{-1}^{1}\left| x \right|dx=\int\limits_{-1}^{0}(-x)dx+\int\limits_{0}^{1}xdx\] can you handle it from here?